OpenStudy (anonymous):

what is the sum of 2y/y+5 and 10/y+5 i need help

OpenStudy (anonymous):

THese have common denominators, so just combine the numerators. \[\frac{2y}{y + 5} + \frac{10}{y + 5}\] \[\frac{2y + 10}{y + 5}\]

OpenStudy (anonymous):

you can simplify that... \[\large \frac{2y+10}{y+5}=\frac{2\cancel{(y+5)}}{\cancel{y+5}}=2 \]

OpenStudy (anonymous):

Oh yeah. I forgot about that.

OpenStudy (anonymous):

thank you guys